OpenStudy (anonymous):

x^4-16^2 find the x intercepts are they 4 and -4 only or is there a 0 intercept

OpenStudy (anonymous):

To find x-intercepts, let y = 0, and solve for x. So you have 0 = x^4 - 16x^2. Can you solve for all x?

OpenStudy (anonymous):

no it would be 0=0 right

OpenStudy (anonymous):

so just 4 and -4 then

OpenStudy (anonymous):

@normman1978 Can you show me your work how you solved?

OpenStudy (anonymous):

@normman1978 I would like to see your work.

OpenStudy (anonymous):

ok my mistake its x^4-16^2=0 x^4=16^2 x^4=256 x=4,-4

OpenStudy (anonymous):

@normman1978 You haven't shown me how you solved the equation. If you dont know, tell me...so I can help. We are here to learn.

OpenStudy (anonymous):



OpenStudy (anonymous):

was that not right

OpenStudy (anonymous):

It hits 0 twice.

OpenStudy (anonymous):

@normman1978 I wanted to see an algebraic solution to the equation. Apparently, you are not responding.

OpenStudy (anonymous):

What I did is called synthetic division

OpenStudy (anonymous):

can you not see my answers

OpenStudy (anonymous):

\[ x^4-16^2=0 x^4=16^2 x^4=256 x=4,-4\]

OpenStudy (anonymous):

@normman1978 I am not interested in your answers if you dont show me your work ...as to how you solved the equation. If you don't want to show me, that is your choice. I just offer to help. Good luck.

OpenStudy (anonymous):

What are you talking about u said have y = 0 and i did \[x^4 -16^2 = 0\]

OpenStudy (anonymous):

\[x^4=-16^2\]

OpenStudy (anonymous):

\[x^4=256\]

OpenStudy (anonymous):

\[\sqrt[4]{x}=\sqrt[4]{256}\]